OpenStudy (aaronandyson):

A bucket is raised from a well by means of a ropw which is wound around a wheel of diameter 77cm.Given that the bucket ascends in 1 minute 28 seconds with a uniform speed of 1.1m/s , calc the number of revolutions the wheel makes in raising the bucket pi = 22/7

OpenStudy (aaronandyson):

@hba

OpenStudy (hba):

Can you like visualize it using the draw tool

OpenStudy (aaronandyson):

n = distance/2*pi*r

OpenStudy (aaronandyson):

88*1.1 = 96.8

OpenStudy (aaronandyson):

we need to convert cm to m

OpenStudy (aaronandyson):

@rvc

OpenStudy (michele_laino):

first step the distance \(d\) covered by the rope linked to the bucket, is: \[\huge d = vt = 1.1 \cdot 88 = ...?\]

OpenStudy (michele_laino):

since 1 minute and 28 seconds are equal to \(88\) seconds

OpenStudy (michele_laino):

@AaronAndyson

OpenStudy (aaronandyson):

96.8

OpenStudy (aaronandyson):

@Michele_Laino

rvc (rvc):

kya huha kya nahi samjha?

OpenStudy (aaronandyson):

pora question

OpenStudy (aaronandyson):

math mera ok ok h

rvc (rvc):

physics hai ye correct

OpenStudy (aaronandyson):

but yeh circles area perimeter wala sums dimag khrb kardete hain

OpenStudy (aaronandyson):

nahi nahi ji this has to do with circumference of a cirlce

rvc (rvc):

haan circumference kya part nahi samjha?

OpenStudy (aaronandyson):

smjha h par yeh revolution wala nhi

rvc (rvc):

what do you mean by revolution?

OpenStudy (aaronandyson):

read the question again @rvc

rvc (rvc):

i know boy will u answer my question :)

OpenStudy (aaronandyson):

first we should write all the given data

OpenStudy (aaronandyson):

n = ? speed = 1.1 ms time = 88s pi = 22/7 circumference = 2*pi*r

OpenStudy (aaronandyson):

diamter of wheel = 77 cm

OpenStudy (aaronandyson):

thats .77 m ryt..??

rvc (rvc):

revolution means one complete cycle in a circle one revolution will be the circumference of the circle correct?

OpenStudy (aaronandyson):

yes

rvc (rvc):

so calculate it :)

OpenStudy (aaronandyson):

we need the no of revolution not the circumference...

rvc (rvc):

1cm = 0.01m

rvc (rvc):

\(\rm N=2 \pi r\\~N=\pi \cdot~d\\~N~is~the~no~of~revolution\)

OpenStudy (aaronandyson):

tsk tsk

rvc (rvc):

tsk?

OpenStudy (aaronandyson):

they have not askedthe circumference yaar

rvc (rvc):

idk im confused

OpenStudy (michele_laino):

second step: the length of the circumference of the wheel is: \[\huge C = 2\pi r \simeq 6.28 \cdot 77 = ...?\]

OpenStudy (michele_laino):

oops.. I have made an error, here is the right formula: \[\huge C = 2\pi r \simeq 6.28 \cdot \frac{{0.77}}{2} = ...?\] please continue

OpenStudy (aaronandyson):

2.14

OpenStudy (michele_laino):

I got: \(C=2.41\)

OpenStudy (michele_laino):

the requested number \(N\) is therefore: \[\huge N = \frac{d}{C} = \frac{{96.8}}{{2.42}} = ...?\]

OpenStudy (aaronandyson):

40.1

OpenStudy (michele_laino):

correct! \(N=40\)

OpenStudy (aaronandyson):

0.77 m is the diameter so should we divide by 2 to get the radius..?

OpenStudy (michele_laino):

yes!

rvc (rvc):

2r=d

OpenStudy (michele_laino):

that's right!! :) @rvc

rvc (rvc):

since the formula has 2 pi r u can club 2r as d then it will be pi d